# A shared compilation stack for distributed-memory parallelism in stencil DSLs

### v1

@@ 3. Sharing Abstractions through IRs
The abstractions transformed by compilers must concretely be realized in code as some IR. IRs that maintain SSA form (Cytron et al., 1991) have proven particularly useful as their direct encoding of a program’s data flow allows transformations without requiring complex analyses. Recently, the MLIR project (Lattner et al., 2021) demonstrated that adding hierarchical structure to SSA-based IRs via nested regions facilitates the modeling of higher-level domain-specific constructs, enabling development of compilers in deep learning (Paszke et al., 2019; Abadi et al., 2016; Abadi et al., 2015), fully-homomorphic encryption (Govindarajan and Moses, 2020), and hardware design (Eldridge et al., 2021; Majumder and Bondhugula, 2024). This SSA+Regions concept provides a principled way to interface distinct domain-specific IRs, which we adopt as a foundational principle in our compiler stack.
@@ 6.2. PSyclone
We have selected two benchmarks to explore performance for our approach with PSyclone, the first is the Piacsek and Williams advection scheme (Piacsek and Williams, 1970), commonly used by Met Office codes such as the MONC high-resolution atmospheric model (Brown et al., 2015), for calculating the movement of quantities through the atmosphere due to kinetic effects (e.g. wind). The second benchmark is the tracer advection kernel from the NEMO ocean model and part of the PSyclone benchmark suite (Science and Technology Facilities Council (STFC), 2021). Both these schemes are extracted from production codes and the benchmarks are different in a couple of crucial ways. Firstly, PW advection contains three separate stencil computations across three fields, whereas tracer advection comprises 24 stencil computations across six fields. Each stencil computation involves many individual calculations, and for tracer advection, there are many more computations than fields due to dependencies between computations generating intermediate results. Furthermore, the tracer advection benchmark contains an outer loop where the entire stencil calculation is executed for 100 iterations.
@@ Acknowledgements.
- Abadi et al. (2015) Martín Abadi, Ashish Agarwal, Paul Barham, Eugene Brevdo, Zhifeng Chen, Craig Citro, Greg S. Corrado, Andy Davis, Jeffrey Dean, Matthieu Devin, Sanjay Ghemawat, Ian Goodfellow, Andrew Harp, Geoffrey Irving, Michael Isard, Yangqing Jia, Rafal Jozefowicz, Lukasz Kaiser, Manjunath Kudlur, Josh Levenberg, Dandelion Mané, Rajat Monga, Sherry Moore, Derek Murray, Chris Olah, Mike Schuster, Jonathon Shlens, Benoit Steiner, Ilya Sutskever, Kunal Talwar, Paul Tucker, Vincent Vanhoucke, Vijay Vasudevan, Fernanda Viégas, Oriol Vinyals, Pete Warden, Martin Wattenberg, Martin Wicke, Yuan Yu, and Xiaoqiang Zheng. 2015. TensorFlow: Large-Scale Machine Learning on Heterogeneous Systems. (2015). https://www.tensorflow.org/ Software available from tensorflow.org.
- Abadi et al. (2016) Martín Abadi, Paul Barham, Jianmin Chen, Zhifeng Chen, Andy Davis, Jeffrey Dean, Matthieu Devin, Sanjay Ghemawat, Geoffrey Irving, Michael Isard, Manjunath Kudlur, Josh Levenberg, Rajat Monga, Sherry Moore, Derek G. Murray, Benoit Steiner, Paul Tucker, Vijay Vasudevan, Pete Warden, Martin Wicke, Yuan Yu, and Xiaoqiang Zheng. 2016. TensorFlow: a system for large-scale machine learning. In Proceedings of the 12th USENIX Conference on Operating Systems Design and Implementation (OSDI’16). USENIX Association, USA, 265–283. https://dl.acm.org/doi/10.5555/3026877.3026899
@@
- Bisbas et al. (2023) George Bisbas, Rhodri Nelson, Mathias Louboutin, Paul H. J. Kelly, Fabio Luporini, and Gerard Gorman. 2023. Automated MPI code generation for scalable finite-difference solvers. (2023). arXiv:cs.DC/2312.13094
@@
- Brown et al. (2015) Nick Brown, Michele Weiland, Adrian Hill, Ben Shipway, Chris Maynard, Thomas Allen, and Mike Rezny. 2015. A highly scalable Met Office NERC Cloud model. In Proceedings of the 3rd International Conference on Exascale Applications and Software (EASC ’15). University of Edinburgh, GBR, 132–137.
@@
- Castillo et al. (2019) Emilio Castillo, Nikhil Jain, Marc Casas, Miquel Moreto, Martin Schulz, Ramon Beivide, Mateo Valero, and Abhinav Bhatele. 2019. Optimizing Computation-Communication Overlap in Asynchronous Task-Based Programs. In Proceedings of the ACM International Conference on Supercomputing (ICS ’19). Association for Computing Machinery, New York, NY, USA, 380–391. https://doi.org/10.1145/3330345.3330379
@@
- Denniston et al. (2016) Tyler Denniston, Shoaib Kamil, and Saman Amarasinghe. 2016. Distributed Halide. In Proceedings of the 21st ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming (PPoPP ’16). Association for Computing Machinery, New York, NY, USA, Article 5, 12 pages. https://doi.org/10.1145/2851141.2851157
@@
- Essadki et al. (2023) Mohamed Essadki, Bertrand Michel, Bruno Maugars, Oleksandr Zinenko, Nicolas Vasilache, and Albert Cohen. 2023. Code Generation for In-Place Stencils. In Proceedings of the 21st ACM/IEEE International Symposium on Code Generation and Optimization (CGO 2023). Association for Computing Machinery, New York, NY, USA, 2–13. https://doi.org/10.1145/3579990.3580006
- Fehr et al. (2023) Mathieu Fehr, Michel Weber, Christian Ulmann, Alexandre Lopoukhine, Martin Lücke, Théo Degioanni, Michel Steuwer, and Tobias Grosser. 2023. Sidekick compilation with xDSL. (2023). arXiv:cs.PL/2311.07422
- Govindarajan and Moses (2020) Sanath Govindarajan and William S. Moses. 2020. SyFER-MLIR: Integrating Fully Homomorphic Encryption Into the MLIR Compiler Framework. (2020). https://math.mit.edu/research/highschool/primes/materials/2020/Govindarajan-Moses.pdf
@@
- Hammond et al. (2023) Jeff Hammond, Lisandro Dalcin, Erik Schnetter, Marc Pérache, Jean-Baptiste Besnard, Jed Brown, Gonzalo Brito Gadeschi, Simon Byrne, Joseph Schuchart, and Hui Zhou. 2023. MPI Application Binary Interface Standardization. In Proceedings of the 30th European MPI Users’ Group Meeting (EuroMPI ’23). Association for Computing Machinery, New York, NY, USA, Article 1, 12 pages. https://doi.org/10.1145/3615318.3615319
@@
- Klöckner (2014) Andreas Klöckner. 2014. Loo.py: transformation-based code generation for GPUs and CPUs. In Proceedings of ACM SIGPLAN International Workshop on Libraries, Languages, and Compilers for Array Programming (ARRAY’14). Association for Computing Machinery, New York, NY, USA, 82–87. https://doi.org/10.1145/2627373.2627387
@@
- Lam et al. (2015) Siu Kwan Lam, Antoine Pitrou, and Stanley Seibert. 2015. Numba: a LLVM-based Python JIT compiler. In Proceedings of the Second Workshop on the LLVM Compiler Infrastructure in HPC (LLVM ’15). Association for Computing Machinery, New York, NY, USA, Article 7, 6 pages. https://doi.org/10.1145/2833157.2833162
@@
- Li et al. (2021) Mingzhen Li, Yi Liu, Hailong Yang, Yongmin Hu, Qingxiao Sun, Bangduo Chen, Xin You, Xiaoyan Liu, Zhongzhi Luan, and Depei Qian. 2021. Automatic Code Generation and Optimization of Large-Scale Stencil Computation on Many-Core Processors. In Proceedings of the 50th International Conference on Parallel Processing (ICPP ’21). ACM, New York, NY, USA, Article 34, 12 pages. https://doi.org/10.1145/3472456.3473517
- Liao et al. (2015) Chunhua Liao, Pei-Hung Lin, Daniel J. Quinlan, Yue Zhao, and Xipeng Shen. 2015. Enhancing Domain Specific Language Implementations through Ontology. In Proceedings of the 5th International Workshop on Domain-Specific Languages and High-Level Frameworks for High Performance Computing (WOLFHPC ’15). ACM, New York, NY, USA, Article 3, 9 pages. https://doi.org/10.1145/2830018.2830022
@@
- Majumder and Bondhugula (2024) Kingshuk Majumder and Uday Bondhugula. 2024. HIR: An MLIR-based Intermediate Representation for Hardware Accelerator Description. (2024), 13 pages. https://doi.org/10.1145/3623278.3624767
- Maruyama et al. (2011) Naoya Maruyama, Tatsuo Nomura, Kento Sato, and Satoshi Matsuoka. 2011. Physis: An Implicitly Parallel Programming Model for Stencil Computations on Large-Scale GPU-Accelerated Supercomputers. In Proceedings of 2011 International Conference for High Performance Computing, Networking, Storage and Analysis (SC ’11). ACM, New York, NY, USA, Article 11, 12 pages. https://doi.org/10.1145/2063384.2063398
@@
- Mitenkov et al. (2023) George Mitenkov, Ioannis Magkanaris, Omar Awile, Pramod Kumbhar, Felix Schürmann, and Alastair F. Donaldson. 2023. MOD2IR: High-Performance Code Generation for a Biophysically Detailed Neuronal Simulation DSL. In Proceedings of the 32nd ACM SIGPLAN International Conference on Compiler Construction (CC 2023). ACM, New York, NY, USA, 203–215. https://doi.org/10.1145/3578360.3580268
- Modular ([n. d.]) Modular. [n. d.]. Mojo: Programming language for all of AI. ([n. d.]). https://www.modular.com/mojo
@@
- Mullapudi et al. (2015) Ravi Teja Mullapudi, Vinay Vasista, and Uday Bondhugula. 2015. PolyMage: Automatic Optimization for Image Processing Pipelines. In Proceedings of the Twentieth International Conference on Architectural Support for Programming Languages and Operating Systems (ASPLOS ’15). Association for Computing Machinery, New York, NY, USA, 429–443. https://doi.org/10.1145/2694344.2694364
- Omlin et al. (2022) Samuel Omlin, Ludovic Räss, and Ivan Utkin. 2022. Distributed Parallelization of xPU Stencil Computations in Julia. (2022). arXiv:cs.DC/2211.15716
@@
- Ravishankar et al. (2015) Mahesh Ravishankar, Justin Holewinski, and Vinod Grover. 2015. Forma: A DSL for Image Processing Applications to Target GPUs and Multi-Core CPUs. In Proceedings of the 8th Workshop on General Purpose Processing Using GPUs (GPGPU-8). Association for Computing Machinery, New York, NY, USA, 109–120. https://doi.org/10.1145/2716282.2716290
- Rawat et al. (2015) Prashant Rawat, Martin Kong, Tom Henretty, Justin Holewinski, Kevin Stock, Louis-Noël Pouchet, J. Ramanujam, Atanas Rountev, and P. Sadayappan. 2015. SDSLc: A Multi-Target Domain-Specific Compiler for Stencil Computations. In Proceedings of the 5th International Workshop on Domain-Specific Languages and High-Level Frameworks for High Performance Computing (WOLFHPC ’15). ACM, New York, NY, USA, Article 6, 10 pages. https://doi.org/10.1145/2830018.2830025
@@
- Rodriguez-Canal et al. (2023b) Gabriel Rodriguez-Canal, Nick Brown, Maurice Jamieson, Emilien Bauer, Anton Lydike, and Tobias Grosser. 2023b. Stencil-HMLS: A multi-layered approach to the automatic optimisation of stencil codes on FPGA. In Proceedings of the SC ’23 Workshops of The International Conference on High Performance Computing, Network, Storage, and Analysis (SC-W ’23). Association for Computing Machinery, New York, NY, USA, 556–565. https://doi.org/10.1145/3624062.3624543
- Science and Technology Facilities Council (STFC) (2021) Science and Technology Facilities Council (STFC). 2021. PSycloneBench: small benchmarks used to inform the development of the PSyclone Domain-Specific Compiler. (2021). https://github.com/stfc/PSycloneBench Accessed 2023-04-24.
- Shajii et al. (2023) Ariya Shajii, Gabriel Ramirez, Haris Smajlović, Jessica Ray, Bonnie Berger, Saman Amarasinghe, and Ibrahim Numanagić. 2023. Codon: A Compiler for High-Performance Pythonic Applications and DSLs. In Proceedings of the 32nd ACM SIGPLAN International Conference on Compiler Construction (CC 2023). ACM, New York, NY, USA, 191–202. https://doi.org/10.1145/3578360.3580275
@@
- Tang et al. (2011) Yuan Tang, Rezaul Alam Chowdhury, Bradley C. Kuszmaul, Chi-Keung Luk, and Charles E. Leiserson. 2011. The Pochoir Stencil Compiler. In Proceedings of the Twenty-third Annual ACM Symposium on Parallelism in Algorithms and Architectures (SPAA ’11). ACM, New York, NY, USA, 117–128. https://doi.org/10.1145/1989493.1989508
- Thaler et al. (2019) Felix Thaler, Stefan Moosbrugger, Carlos Osuna, Mauro Bianco, Hannes Vogt, Anton Afanasyev, Lukas Mosimann, Oliver Fuhrer, Thomas C. Schulthess, and Torsten Hoefler. 2019. Porting the COSMO Weather Model to Manycore CPUs. In Proceedings of the Platform for Advanced Scientific Computing Conference (PASC ’19). Association for Computing Machinery, New York, NY, USA, Article 13, 11 pages. https://doi.org/10.1145/3324989.3325723
@@
- Unat et al. (2011) Didem Unat, Xing Cai, and Scott B. Baden. 2011. Mint: Realizing CUDA Performance in 3D Stencil Methods with Annotated C. In Proceedings of the International Conference on Supercomputing (ICS ’11). ACM, New York, NY, USA, 214–224. https://doi.org/10.1145/1995896.1995932
@@
- Zhao et al. (2021) Tuowen Zhao, Mary Hall, Hans Johansen, and Samuel Williams. 2021. Improving Communication by Optimizing On-Node Data Movement with Data Layout. In Proceedings of the 26th ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming (PPoPP ’21). Association for Computing Machinery, New York, NY, USA, 304–317. https://doi.org/10.1145/3437801.3441598
@@
- Ziogas et al. (2021) Alexandros Nikolaos Ziogas, Timo Schneider, Tal Ben-Nun, Alexandru Calotoiu, Tiziano De Matteis, Johannes de Fine Licht, Luca Lavarini, and Torsten Hoefler. 2021. Productivity, Portability, Performance: Data-Centric Python. In Proceedings of the International Conference for High Performance Computing, Networking, Storage and Analysis (SC ’21). Association for Computing Machinery, New York, NY, USA, Article 95, 13 pages. https://doi.org/10.1145/3458817.3476176

### v2

@@ 3. Sharing Abstractions through IRs
The abstractions transformed by compilers must concretely be realized in code as some IR. IRs that maintain SSA form (Cytron et al., 1991) have proven particularly useful as their direct encoding of a program’s data flow allows transformations without requiring complex analyses. Recently, the MLIR project (Lattner et al., 2021) demonstrated that adding hierarchical structure to SSA-based IRs via nested regions facilitates the modeling of higher-level domain-specific constructs, enabling development of compilers in deep learning (Paszke et al., 2019; Abadi et al., 2016, 2015), fully-homomorphic encryption (Govindarajan and Moses, 2020), and hardware design (Eldridge et al., 2021; Majumder and Bondhugula, 2024). This SSA+Regions concept provides a principled way to interface distinct domain-specific IRs, which we adopt as a foundational principle in our compiler stack.
@@ 6.2. PSyclone
We have selected two benchmarks to explore performance for our approach with PSyclone, the first is the Piacsek and Williams advection scheme (Piacsek and Williams, 1970), commonly used by Met Office codes such as the MONC high-resolution atmospheric model (Brown et al., 2015), for calculating the movement of quantities through the atmosphere due to kinetic effects (e.g. wind). The second benchmark is the tracer advection kernel from the NEMO ocean model and part of the PSyclone benchmark suite (Science and Technology Facilities Council (2021), STFC). Both these schemes are extracted from production codes and the benchmarks are different in a couple of crucial ways. Firstly, PW advection contains three separate stencil computations across three fields, whereas tracer advection comprises 24 stencil computations across six fields. Each stencil computation involves many individual calculations, and for tracer advection, there are many more computations than fields due to dependencies between computations generating intermediate results. Furthermore, the tracer advection benchmark contains an outer loop where the entire stencil calculation is executed for 100 iterations.
@@ Acknowledgements.
- Abadi et al. (2015) Martín Abadi, Ashish Agarwal, Paul Barham, Eugene Brevdo, Zhifeng Chen, Craig Citro, Greg S. Corrado, Andy Davis, Jeffrey Dean, Matthieu Devin, Sanjay Ghemawat, Ian Goodfellow, Andrew Harp, Geoffrey Irving, Michael Isard, Yangqing Jia, Rafal Jozefowicz, Lukasz Kaiser, Manjunath Kudlur, Josh Levenberg, Dandelion Mané, Rajat Monga, Sherry Moore, Derek Murray, Chris Olah, Mike Schuster, Jonathon Shlens, Benoit Steiner, Ilya Sutskever, Kunal Talwar, Paul Tucker, Vincent Vanhoucke, Vijay Vasudevan, Fernanda Viégas, Oriol Vinyals, Pete Warden, Martin Wattenberg, Martin Wicke, Yuan Yu, and Xiaoqiang Zheng. 2015. TensorFlow: Large-Scale Machine Learning on Heterogeneous Systems. https://www.tensorflow.org/ Software available from tensorflow.org.
- Abadi et al. (2016) Martín Abadi, Paul Barham, Jianmin Chen, Zhifeng Chen, Andy Davis, Jeffrey Dean, Matthieu Devin, Sanjay Ghemawat, Geoffrey Irving, Michael Isard, Manjunath Kudlur, Josh Levenberg, Rajat Monga, Sherry Moore, Derek G. Murray, Benoit Steiner, Paul Tucker, Vijay Vasudevan, Pete Warden, Martin Wicke, Yuan Yu, and Xiaoqiang Zheng. 2016. TensorFlow: a system for large-scale machine learning. In Proceedings of the 12th USENIX Conference on Operating Systems Design and Implementation (Savannah, GA, USA) (OSDI’16). USENIX Association, USA, 265–283. https://dl.acm.org/doi/10.5555/3026877.3026899
@@
- Bisbas et al. (2023) George Bisbas, Rhodri Nelson, Mathias Louboutin, Paul H. J. Kelly, Fabio Luporini, and Gerard Gorman. 2023. Automated MPI code generation for scalable finite-difference solvers. arXiv:2312.13094 [cs.DC]
@@
- Brown et al. (2015) Nick Brown, Michele Weiland, Adrian Hill, Ben Shipway, Chris Maynard, Thomas Allen, and Mike Rezny. 2015. A highly scalable Met Office NERC Cloud model. In Proceedings of the 3rd International Conference on Exascale Applications and Software (Edinburgh, UK) (EASC ’15). University of Edinburgh, GBR, 132–137.
@@
- Castillo et al. (2019) Emilio Castillo, Nikhil Jain, Marc Casas, Miquel Moreto, Martin Schulz, Ramon Beivide, Mateo Valero, and Abhinav Bhatele. 2019. Optimizing Computation-Communication Overlap in Asynchronous Task-Based Programs. In Proceedings of the ACM International Conference on Supercomputing (Phoenix, Arizona) (ICS ’19). Association for Computing Machinery, New York, NY, USA, 380–391. https://doi.org/10.1145/3330345.3330379
@@
- Denniston et al. (2016) Tyler Denniston, Shoaib Kamil, and Saman Amarasinghe. 2016. Distributed Halide. In Proceedings of the 21st ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming (Barcelona, Spain) (PPoPP ’16). Association for Computing Machinery, New York, NY, USA, Article 5, 12 pages. https://doi.org/10.1145/2851141.2851157
@@
- Essadki et al. (2023) Mohamed Essadki, Bertrand Michel, Bruno Maugars, Oleksandr Zinenko, Nicolas Vasilache, and Albert Cohen. 2023. Code Generation for In-Place Stencils. In Proceedings of the 21st ACM/IEEE International Symposium on Code Generation and Optimization (Montréal, QC, Canada) (CGO 2023). Association for Computing Machinery, New York, NY, USA, 2–13. https://doi.org/10.1145/3579990.3580006
- Fehr et al. (2023) Mathieu Fehr, Michel Weber, Christian Ulmann, Alexandre Lopoukhine, Martin Lücke, Théo Degioanni, Michel Steuwer, and Tobias Grosser. 2023. Sidekick compilation with xDSL. arXiv:2311.07422 [cs.PL]
- Govindarajan and Moses (2020) Sanath Govindarajan and William S. Moses. 2020. SyFER-MLIR: Integrating Fully Homomorphic Encryption Into the MLIR Compiler Framework. https://math.mit.edu/research/highschool/primes/materials/2020/Govindarajan-Moses.pdf
@@
- Hammond et al. (2023) Jeff Hammond, Lisandro Dalcin, Erik Schnetter, Marc Pérache, Jean-Baptiste Besnard, Jed Brown, Gonzalo Brito Gadeschi, Simon Byrne, Joseph Schuchart, and Hui Zhou. 2023. MPI Application Binary Interface Standardization. In Proceedings of the 30th European MPI Users’ Group Meeting (<conf-loc>, <city>Bristol</city>, <country>United Kingdom</country>, </conf-loc>) (EuroMPI ’23). Association for Computing Machinery, New York, NY, USA, Article 1, 12 pages. https://doi.org/10.1145/3615318.3615319
@@
- Klöckner (2014) Andreas Klöckner. 2014. Loo.py: transformation-based code generation for GPUs and CPUs. In Proceedings of ACM SIGPLAN International Workshop on Libraries, Languages, and Compilers for Array Programming (Edinburgh, United Kingdom) (ARRAY’14). Association for Computing Machinery, New York, NY, USA, 82–87. https://doi.org/10.1145/2627373.2627387
@@
- Lam et al. (2015) Siu Kwan Lam, Antoine Pitrou, and Stanley Seibert. 2015. Numba: a LLVM-based Python JIT compiler. In Proceedings of the Second Workshop on the LLVM Compiler Infrastructure in HPC (Austin, Texas) (LLVM ’15). Association for Computing Machinery, New York, NY, USA, Article 7, 6 pages. https://doi.org/10.1145/2833157.2833162
@@
- Li et al. (2021) Mingzhen Li, Yi Liu, Hailong Yang, Yongmin Hu, Qingxiao Sun, Bangduo Chen, Xin You, Xiaoyan Liu, Zhongzhi Luan, and Depei Qian. 2021. Automatic Code Generation and Optimization of Large-Scale Stencil Computation on Many-Core Processors. In Proceedings of the 50th International Conference on Parallel Processing (Lemont, IL, USA) (ICPP ’21). ACM, New York, NY, USA, Article 34, 12 pages. https://doi.org/10.1145/3472456.3473517
- Liao et al. (2015) Chunhua Liao, Pei-Hung Lin, Daniel J. Quinlan, Yue Zhao, and Xipeng Shen. 2015. Enhancing Domain Specific Language Implementations through Ontology. In Proceedings of the 5th International Workshop on Domain-Specific Languages and High-Level Frameworks for High Performance Computing (Austin, Texas) (WOLFHPC ’15). ACM, New York, NY, USA, Article 3, 9 pages. https://doi.org/10.1145/2830018.2830022
@@
- Majumder and Bondhugula (2024) Kingshuk Majumder and Uday Bondhugula. 2024. HIR: An MLIR-based Intermediate Representation for Hardware Accelerator Description. , 13 pages. https://doi.org/10.1145/3623278.3624767
- Maruyama et al. (2011) Naoya Maruyama, Tatsuo Nomura, Kento Sato, and Satoshi Matsuoka. 2011. Physis: An Implicitly Parallel Programming Model for Stencil Computations on Large-Scale GPU-Accelerated Supercomputers. In Proceedings of 2011 International Conference for High Performance Computing, Networking, Storage and Analysis (Seattle, Washington) (SC ’11). ACM, New York, NY, USA, Article 11, 12 pages. https://doi.org/10.1145/2063384.2063398
@@
- Mitenkov et al. (2023) George Mitenkov, Ioannis Magkanaris, Omar Awile, Pramod Kumbhar, Felix Schürmann, and Alastair F. Donaldson. 2023. MOD2IR: High-Performance Code Generation for a Biophysically Detailed Neuronal Simulation DSL. In Proceedings of the 32nd ACM SIGPLAN International Conference on Compiler Construction (Montréal, QC, Canada) (CC 2023). ACM, New York, NY, USA, 203–215. https://doi.org/10.1145/3578360.3580268
- Modular ([n. d.]) Modular. [n. d.]. Mojo: Programming language for all of AI. https://www.modular.com/mojo
@@
- Mullapudi et al. (2015) Ravi Teja Mullapudi, Vinay Vasista, and Uday Bondhugula. 2015. PolyMage: Automatic Optimization for Image Processing Pipelines. In Proceedings of the Twentieth International Conference on Architectural Support for Programming Languages and Operating Systems (Istanbul, Turkey) (ASPLOS ’15). Association for Computing Machinery, New York, NY, USA, 429–443. https://doi.org/10.1145/2694344.2694364
- Omlin et al. (2022) Samuel Omlin, Ludovic Räss, and Ivan Utkin. 2022. Distributed Parallelization of xPU Stencil Computations in Julia. arXiv:2211.15716 [cs.DC]
@@
- Ravishankar et al. (2015) Mahesh Ravishankar, Justin Holewinski, and Vinod Grover. 2015. Forma: A DSL for Image Processing Applications to Target GPUs and Multi-Core CPUs. In Proceedings of the 8th Workshop on General Purpose Processing Using GPUs (San Francisco, CA, USA) (GPGPU-8). Association for Computing Machinery, New York, NY, USA, 109–120. https://doi.org/10.1145/2716282.2716290
- Rawat et al. (2015) Prashant Rawat, Martin Kong, Tom Henretty, Justin Holewinski, Kevin Stock, Louis-Noël Pouchet, J. Ramanujam, Atanas Rountev, and P. Sadayappan. 2015. SDSLc: A Multi-Target Domain-Specific Compiler for Stencil Computations. In Proceedings of the 5th International Workshop on Domain-Specific Languages and High-Level Frameworks for High Performance Computing (Austin, Texas) (WOLFHPC ’15). ACM, New York, NY, USA, Article 6, 10 pages. https://doi.org/10.1145/2830018.2830025
@@
- Rodriguez-Canal et al. (2023b) Gabriel Rodriguez-Canal, Nick Brown, Maurice Jamieson, Emilien Bauer, Anton Lydike, and Tobias Grosser. 2023b. Stencil-HMLS: A multi-layered approach to the automatic optimisation of stencil codes on FPGA. In Proceedings of the SC ’23 Workshops of The International Conference on High Performance Computing, Network, Storage, and Analysis (<conf-loc>, <city>Denver</city>, <state>CO</state>, <country>USA</country>, </conf-loc>) (SC-W ’23). Association for Computing Machinery, New York, NY, USA, 556–565. https://doi.org/10.1145/3624062.3624543
- Science and Technology Facilities Council (2021) (STFC) Science and Technology Facilities Council (STFC). 2021. PSycloneBench: small benchmarks used to inform the development of the PSyclone Domain-Specific Compiler. https://github.com/stfc/PSycloneBench Accessed 2023-04-24.
- Shajii et al. (2023) Ariya Shajii, Gabriel Ramirez, Haris Smajlović, Jessica Ray, Bonnie Berger, Saman Amarasinghe, and Ibrahim Numanagić. 2023. Codon: A Compiler for High-Performance Pythonic Applications and DSLs. In Proceedings of the 32nd ACM SIGPLAN International Conference on Compiler Construction (Montréal, QC, Canada) (CC 2023). ACM, New York, NY, USA, 191–202. https://doi.org/10.1145/3578360.3580275
@@
- Tang et al. (2011) Yuan Tang, Rezaul Alam Chowdhury, Bradley C. Kuszmaul, Chi-Keung Luk, and Charles E. Leiserson. 2011. The Pochoir Stencil Compiler. In Proceedings of the Twenty-third Annual ACM Symposium on Parallelism in Algorithms and Architectures (San Jose, California, USA) (SPAA ’11). ACM, New York, NY, USA, 117–128. https://doi.org/10.1145/1989493.1989508
- Thaler et al. (2019) Felix Thaler, Stefan Moosbrugger, Carlos Osuna, Mauro Bianco, Hannes Vogt, Anton Afanasyev, Lukas Mosimann, Oliver Fuhrer, Thomas C. Schulthess, and Torsten Hoefler. 2019. Porting the COSMO Weather Model to Manycore CPUs. In Proceedings of the Platform for Advanced Scientific Computing Conference (Zurich, Switzerland) (PASC ’19). Association for Computing Machinery, New York, NY, USA, Article 13, 11 pages. https://doi.org/10.1145/3324989.3325723
@@
- Unat et al. (2011) Didem Unat, Xing Cai, and Scott B. Baden. 2011. Mint: Realizing CUDA Performance in 3D Stencil Methods with Annotated C. In Proceedings of the International Conference on Supercomputing (Tucson, Arizona, USA) (ICS ’11). ACM, New York, NY, USA, 214–224. https://doi.org/10.1145/1995896.1995932
@@
- Zhao et al. (2021) Tuowen Zhao, Mary Hall, Hans Johansen, and Samuel Williams. 2021. Improving Communication by Optimizing On-Node Data Movement with Data Layout. In Proceedings of the 26th ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming (Virtual Event, Republic of Korea) (PPoPP ’21). Association for Computing Machinery, New York, NY, USA, 304–317. https://doi.org/10.1145/3437801.3441598
@@
- Ziogas et al. (2021) Alexandros Nikolaos Ziogas, Timo Schneider, Tal Ben-Nun, Alexandru Calotoiu, Tiziano De Matteis, Johannes de Fine Licht, Luca Lavarini, and Torsten Hoefler. 2021. Productivity, Portability, Performance: Data-Centric Python. In Proceedings of the International Conference for High Performance Computing, Networking, Storage and Analysis (St. Louis, Missouri) (SC ’21). Association for Computing Machinery, New York, NY, USA, Article 95, 13 pages. https://doi.org/10.1145/3458817.3476176